OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is (4x8 + 3x6) cubic feet. What polynomial represents the height of the box in feet

OpenStudy (anonymous):

\[V=lwh\] \[4x^8+3x^6=x * x * h\] \[4x^8+3x^6=x^2h\] \[h=\frac{ 4x^8+3x^6 }{ x^2 }\] \[h=\frac{ x^2(4x^6+3x^4) }{ x^2 }\] the x^2 cancels out and you get \[h=4x^6+ 3x^4\]

OpenStudy (anonymous):

V=Volume l=length w=width h=height